HOW TO FIND THE RANK OF THE WORD "PRISON" IF THE GIVEN WORD IS "SIPRON"

Grade:12

1 Answers

Sukhendra Reddy Rompally B.Tech Mining Machinery Engg, ISM Dhanbad
13 years ago
RANK of a word is its position in the dictionary of words formed using all the letters arranged in all possible ways. Here, using the letters S I P R O N,form the possible words.since,i,n,o come before p in a dictionary,use these as first letters and form words,which is 3*5! .now,using P as first letter,the 2nd letter to be used,which comes before R are I,N,O,which means total words=3*4! The 3rd letter is I,which gets the 1st place in the dictionary.4th letter S,u have options of O and N before.so, 2*2! Now the 5th letter can be N or O,since N comes first,1*1! .the rank=sum of all these + 1,which is the position of the required word.the answer =360+72+4+1+1=438 hope,u r clear with the solution.if not,feel free to call me on 07209736303 PLZZ APPROVE MY ANSWER IF YOU LIKE IT